If m=cosec theta -sin theta and n=sec theta -cos theta,then m^2/3+n^2/3=

Grade:11

1 Answers

Maharshi Shukla
6 years ago
-sintheta×cosec theta =-1 and same way-cos theta× sec theta = -1(-1)^2/3+(-1)^2/3=(-1)^2^1/3+(-1)^2^1/3=1^1/3+1^1/3=1+1=2